A monolayer containing 3.15×10^-6 g of oleic acid has an area of 20.0 cm^2. The density of oleic acid is 0.895 g / mL. What is the thickness of the monolayer (the length of an oleic acid molecule)?
Answer #1

mass   = volume * density

3.15*10-6 = volume *0.895

volume   = 3.15*10-6/0.895   = 3.52*10-6 ml = 3.52*10-6 cm3

volume   = area * thickness

3.52*10-6 = 20*thickness

thickness   = 3.52*10-6/20   = 1.76*10-7 cm >>>> answer